OpenStudy (lilai3):

Hello. Bobby, Billy, and Barry Smith are each one year apart in age. The sum of their ages is greater than the age of their father, who is 60. What is the youngest age that the oldest brother can be? If you help, I will give you free imaginary oreos! :D Y ou O bviously L ove O reos Thank you!~

OpenStudy (the_fizicx99):

Well for oreos sure

OpenStudy (lilai3):

lol yeah.

hartnn (hartnn):

so let the youngest brother be 'x' years old other 2 brothers will be x+1 and x+2 years old. got this ? oldest brother will be x+2 years old

OpenStudy (lilai3):

Okay, but we are still solving for the oldest brother, right?

hartnn (hartnn):

we'll solve for x and for oldest brother, we'll calculate x+2

OpenStudy (lilai3):

\[35\ge15+2h?\]

hartnn (hartnn):

how u got that ?

OpenStudy (lilai3):

Oops.

OpenStudy (lilai3):

a>21 ? Sorry, at first I typed the wrong answer down. LOL.

hartnn (hartnn):

x+x+1+x+2 > 60 x> 19 so, eldest brother = x+2 = a so, x+2>21 a>21 YES< thats correct :)

OpenStudy (lilai3):

On the problem, it stated that the "sum of their ages is greater than the age of their father, who is 60." However, if the oldest brother is 21, that means the other brothers are 20 and 19, correct? Then you add them altogether, and you get, well, I got, 60. But the problem stated that it was" greater than the age of their father, who is 60."

hartnn (hartnn):

a>21 means a can be 22 so take eldest brother age as 22

hartnn (hartnn):

a>21 means a CANNOT be 21 thats why we take a =22

OpenStudy (lilai3):

Then the answer would be 22, wouldn't it?

hartnn (hartnn):

absolutely :)

OpenStudy (lilai3):

Oh, thank you. I get it now. *gives an imaginary Oreo, cuz who won't RESIST them!!!*

hartnn (hartnn):

what is Oreo ?

OpenStudy (the_fizicx99):

:O wuuuut ^

OpenStudy (lilai3):

don't tell me. it's like a kkkooocccieee

hartnn (hartnn):

ohh...just googled it. i never knew about that

OpenStudy (the_fizicx99):

O.e

OpenStudy (lilai3):

oh...wow....now you know!! lol what part of the world are you from?

hartnn (hartnn):

lol, i am from India

OpenStudy (the_fizicx99):

@hartnn ♥

OpenStudy (lilai3):

oh. you should try an Oreo when you get the chance to come to the USA. lol @tHe_FiZiCx99

hartnn (hartnn):

maybe there are Oreos in India too....i'll search the supermarket :)

OpenStudy (lilai3):

lol.

hartnn (hartnn):

by that time, i'll be satisfied with your imaginary one ;)

OpenStudy (the_fizicx99):

dafuqqq ;~; *gives hartnn money for food* take it! :*

OpenStudy (lilai3):

Jamal works in a city and sometimes takes a taxi to work. The taxicabs charge $1.50 for the first 1/5 mile and $0.25 for each additional 1/5 mile. Jamal has only $3.75 in his pocket. What is the maximum distance he can travel by cab if he does not intend to tip the cab driver?

hartnn (hartnn):

let jamal can go 'x' additional 1/5 miles so he will be charged 0.25x dollars for that so, 1.5 + 0.25x is the total charge

OpenStudy (lilai3):

\[1.50+0.25x \le3.75\]

hartnn (hartnn):

but he only have max of 3.75$ so, our equation will be 1.5 +0.25x <= 3.75

OpenStudy (lilai3):

right? ^ ^ ^

hartnn (hartnn):

YES

OpenStudy (lilai3):

oh...wha- what?

OpenStudy (lilai3):

wait..what? i'm confused..what? am i right or wrong? lol

hartnn (hartnn):

YES, means yes, you are right :)

OpenStudy (lilai3):

oh (;

hartnn (hartnn):

so what u get for x ?

OpenStudy (lilai3):

The answer for \[x\] that I got was 2 miles...

hartnn (hartnn):

how ?

OpenStudy (lilai3):

However, I also got 9 miles....

hartnn (hartnn):

x= 9 right ?

OpenStudy (lilai3):

i don't know. i think i got both 2 and 9, just dunno how. ditto.

OpenStudy (lilai3):

and then i got another answer:10.

OpenStudy (lilai3):

is this the three-answer problem or what? :O

hartnn (hartnn):

1.5 +0.25x <= 3.75 0.25x <= 2.25 x <= 9 so he can go 9 (1/5) additional miles which is 9/5 plus 1/5 miles initially so, 9/5 + 1/5 = 10/5 = 2 miles only

OpenStudy (lilai3):

so is it 9 or 2 or 10?

hartnn (hartnn):

2

OpenStudy (lilai3):

...how?

OpenStudy (lilai3):

how?

OpenStudy (lilai3):

THANK YOU AMIGO. THANK YOU. *gives another imaginary Oreo and a medal, which is iconislistic)

OpenStudy (lilai3):

bye. thanks. don't reply thanks bye.